OpenStudy (anonymous):

A particle moves along the x-axis so that at any time t, measured in seconds, its position is given by s(t) = sin(t) − 4cos(2t), measured in feet. What is the acceleration of the particle at time t = π seconds? Numerical Answers Expected!

OpenStudy (jhannybean):

\[\ \int s(t) dt = v(t)+c\]\[\ \int [v(t) + c]dt = a(t) +ct +d\]

OpenStudy (anonymous):

So what do I get as the answer ? :)

OpenStudy (jhannybean):

So you have to integrate the position function twice to get the acceleration function when t = \(\ \pi\)

OpenStudy (jhannybean):

can you integrate the position function and tell me what you get?

OpenStudy (anonymous):

I'm not sure how I'm in a rush so you can give me the answer and then explain it because I have a limited amount of time to submit the work :P Thanks :)

OpenStudy (jhannybean):

Is this a test you are working on, or homework?

OpenStudy (anonymous):

pop quiz

OpenStudy (jhannybean):

It's against Openstudy code of conduct for me to give you answers. I am sorry.

OpenStudy (anonymous):

okay then can u please explain it as fast as u can ?

OpenStudy (jhannybean):

If you work with me on the problem then I can help you :)

OpenStudy (anonymous):

ok :) how do i start

OpenStudy (jhannybean):

Yeah, you integrate both functions to get the function of acceleration, then plug in t = pi, solve for your constant (i.e c, d, e ,f, etc...) and once you get the value of your constant, plug it back into the equation for acceleration and you're set.

OpenStudy (anonymous):

so what do i get as a result ?: )

OpenStudy (anonymous):

i got the answer as 0